OpenStudy (anonymous):

what is the similarity ratio of these figures along with the ratio of their surface area and volume? ( the dimensions of the two similar cubes are 2,2,2 and 4,4,4)

OpenStudy (anonymous):

ratio of sides = 2 : 4 = 1:2 .. areas = 2^2 : 4^2 = 4: 16 = 1 : 2 .. volume = 2^3 : 4^3 = 8 : 64 = 1 : 8

OpenStudy (anonymous):

* oops ratio of areas = 1 : 4 ( not 1:2 )

OpenStudy (anonymous):

thanks!! =) oh and i love your picture btw lol